OpenStudy (vijay222ms):

If sin^2 +sinx = 1 then cos^10 x+5cos^12 x+10cos^14 x+10cos^16 x+5cos^18 x+cos^20 x

OpenStudy (vijay222ms):

@Compassionate @dan815 @mathslover @nincompoop @TheSmartOne @eliassaab @bibby @shrutipande9

OpenStudy (vijay222ms):

@AlexandervonHumboldt2

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\sin^2x+\sin x=1\hspace{.33em}\\~\\ \sin x=1-\sin^2x----\color{red}{(1)}\hspace{.33em}\\~\\ \cos^{10}x+5\cos^{12}x+10\cos^{14}x+10\cos^{16}x+5\cos^{18}x+\cos^{20}x\hspace{.33em}\\~\\ =\cos^{10}x(1+5\cos^{2}x+10\cos^{4}x+10\cos^{6}x+5\cos^{8}x+\cos^{10}x)\hspace{.33em}\\~\\ =\cos^{10}x(1+5\cos^{2}x+10\cos^{4}x+10\cos^{6}x+5\cos^{8}x+\cos^{10}x)\hspace{.33em}\\~\\ \normalsize \text{put} ~~\cos^{2}x=a \hspace{.33em}\\~\\ =a^{5}(1+5a+10a^{2}+10a^{3}+5a^{4}+a^{5})\hspace{.33em}\\~\\ \normalsize \text{this is the binomial expansion of } ~~(a+1)^5 \hspace{.33em}\\~\\ =a^{5}(a+1)^5 \hspace{.33em}\\~\\ =\left(\cos^{2}x(\cos^{2}x+1)\right )^5 \hspace{.33em}\\~\\ =\left(\cos^{4}x+\cos^{2}x\right )^5 \hspace{.33em}\\~\\ =\left((\cos^{2}x)^2+\cos^{2}x\right )^5 \hspace{.33em}\\~\\ =\left((1-\sin^2x)^2+\cos^{2}x\right )^5 \hspace{.33em}\\~\\ =\left(\sin x)^2+\cos^{2}x\right )^5 \hspace{.33em}\\~\\ =1^5 }\end{align}\)

OpenStudy (vijay222ms):

@mathmath333 Thanks